Two jets are in New Orleans. One jet is leaving at 45 mph. The other is leaving at 90 mph in the opposite direction. How many hours will it be before they are 20 miles apart? I have already made the chart and tried using my textbook but the equations in the textbook are different from this world problem. Please help.
Answer by stanbon(75887)

Two jets are in New Orleans. One jet is leaving at 45 mph. The other is leaving at 90 mph in the opposite direction. How many hours will it be before they are 20 miles apart?
------------------
The jets are separating at the rate of (45+90) = 135 mph
----
Equation:
time = distance/rate
time = 20/135 = 4/27 of an hour
---
time = (4/27)60 min = 8.89 minutes
=======================================
Cheers,
Stan H.
=================